OpenStudy (lidia.12):

Let f(x) =3x, g(x)=x+1. Find the composite function. (F•g)(2)

OpenStudy (mhchen):

Same thing as other question you had. This time it's f(g(x)) So first you find g(2) = 2 + 1 = 3 Then plug that inside f(x): f(3) = 3*3 = (?) <-- Solve that part

OpenStudy (turbostudier):

wouldnt it just be: (3x*x+1)(2) ???

OpenStudy (3mar):

@turbostudier No, it would not and for mhchen It is not quite right! The correct procedure is to replace every x in f(x) with all of g(x) That means: \[f(g(x))=3.g(x)=3(x+1)=3x+3\] then substitute with 2 \[f(g(x))=3(2)+3=6+3=9\]